# Hilbert schemes on blowing ups and the free Boson

### v2

@@ 2.2. Highest representation theory of $\mathcal{E}$ and Fermionic Fock space
The verma module of the highest weight representation of weight $\lambda$ is denoted as $F_{\lambda,\mathcal{E}}$, and the hightest weight vector is denoted as 00$\rangle$0 ⟩. We denote $F_{\mathcal{E}}:=F_{1,\mathcal{E}}$.
@@
$E\to\lambda E,\quad K_{1}\to\lambda K_{1},\quad F\to\lambda F,\quad K_{-1}\to\lambda K_{-1},$
and under this auto equivalence, the representation $F_{\lambda,\mathcal{E}}$ is mapping to $F_{\mathcal{E}}$.
Now we explain that $F_{\mathcal{E}}$ is a Fermionic Fock space: given $i\in\mathbb{Z}_{>0}$, we denote $x_{i}$ as a formal symbol. We consider the infinite-dimensional vector space
$V=\bigoplus_{i=1}^{\infty}\mathbb{Q}x_{i}.$
Definition 2.6.
The Fermionic Fock space $F$ is defined as the infinite wedge product
$F:=\bigoplus_{i=1}^{\infty}\wedge^{i}F$
and it has basis
@@
for any sequence $0<i_{1}<i_{2}<\cdots<i_{k},k\geq 0$.
Through straightforward calculations, we find that
@@
There is a representation of $\mathcal{E}$ on the Fermionic Fock space $F$, such that the action of $id$ is identity and
@@
where $x_{0}:=0$. Under the action of $\mathcal{E}$ on $F$, the action of $P_{i}$ is wedging $x_{i}$ and the action of $Q_{i}$ is $\frac{\partial}{\partial x_{i}}$.
@@
Let $|0\rangle$ be the highest weight vector of $F_{\mathcal{E}}$. Then themap
$|0\rangle\to 1$
@@
Now we prove that the Fermionic Fock space $F$ is irreducible as a $\mathcal{E}$-module. Let $a=\oplus_{i=1}^{N}a_{i}\in F$, we take a monomial
@@ 2.3. Admissible representations and Boson-Fermion correspondence
We consider the category of graded $\mathcal{E}$-modules (resp. $\mathcal{C}$-repsentations), i.e. $\mathcal{E}$-modules (resp. $\mathcal{C}$-modules)
@@
A graded $\mathcal{E}$-module
@@
is called admissible if
  1. (1)
    $K_{1}|_{W_{0}}=K_{-1}|_{W_{0}}=id$;
  2. (2)
    for any $x\in W_{i}$ such that $i>0$, there exists $N>0$ such that $K_{-1}^{N}x=0$.
@@
    for any $x\in W$, there exists $N>0$ such that $Q_{m}x=0$ when $m>0$.
@@
On the Fermionic Fock space, we assoiciate $x_{i}$ with degree $1$. Then $F$ is a graded admissible $\mathcal{E}$-module.
@@
$\begin{cases}H_{k}\to y_{k},\quad k>0,\\ H_{k}\to\frac{\partial}{\partial y_{-k}},\quad k<0\end{cases}$
which is an admissible representation.
@@
An $\mathcal{H}$-representation $W$ is called admissible, if for any $x\in W$ and $i>0$, there exists an integer $N>0$ such that $H_{-\lambda}x=0$ for any $|\lambda|>N$.
@@
We have an equivalence among the abelian category of admissible $\mathcal{H}$-modules, graded admissible $\mathcal{C}$-modules and graded admissible $\mathcal{E}$-modules.
@@ 2.4. Graded $\mathcal{E}$-modules
We only need to prove that
@@
First we notice that for any non-zero $x\in\oplus_{i>0}W_{i}$, there exists $M\geq 0$ such $F_{M}x\neq 0$. Otherwise, by 2.3, for any $M>0$
@@
which contradicts with the assumption of admissibility. Let $M_{0}$ be the smallest integer such that $F_{M_{0}}x\neq 0$. Then by induction, we have
$Q_{M+1}x=K_{1}Q_{M}K_{-1}x=\cdots=K_{1}^{M_{0}}Q_{1}K_{-1}^{M_{0}}x=K_{1}^{M_{0}+1}F_{M}x.$
@@
Let $W$ be an admissible $\mathcal{C}$-module. Let
For any operator $T$ on $W$, we consider
$\Gamma(T)(z):=\sum_{i\in\mathbb{Z}}\Gamma_{i}(T)z^{i}$
such that
which is well defined due to the admissible property. We also formally written
$\Gamma(T)(z):=P(z)TQ(z^{-1}).$
We denote
$\Lambda:=\Gamma_{0}(id).$
by Lemma A.4 we have $\Lambda|_{W_{i}}=i\cdot id$ and hence is a invertible operator.
  1. (1)
    $K_{1},K_{-1}$ are identity on $W_{0}$;
  2. (2)
    we have
    $\Lambda K_{1}=\Gamma_{1}(K_{1}),\quad\Lambda K_{-1}=\Gamma_{-1}(K_{-1}).$
We also define
$E:=P_{1}K_{1},\quad F:=K_{-1}Q_{1}.$
In Section A.2, we will prove that
Proposition 2.16.
The operators $K_{1},K_{-1},E,F$ induce an admissible $\mathcal{E}$-module on $W$.
@@ 2.5. Filtration and Fermionization
It satisies that $W_{i}\subset W_{j}$ for $i\leq j$ and
@@
(2.5) $(Id-H_{k}\mathcal{L}_{k})W_{k}\subset W_{k-1}$
and thus the following operators
$E=\bigoplus_{i\geq 0}id,\quad F=\bigoplus_{i\geq 0}(id-H_{k}\mathcal{L}_{k})$
are well defined on
$W=\bigoplus_{i\geq 0}W_{i}.$
Moreover, considering the operators
@@
where $H_{0}=\mathcal{L}_{0}:=id$. Then $W$ is an graded admissible $\mathcal{E}$-module.
@@ 2.6. Stable limit and Bosonization
we define the stable limit, denotes $W_{\infty}$ as the limit
$\lim_{n\to\infty}W_{i}$
where the morphism form $W_{i}$ to $W_{i+1}$ is $K_{1}$ which is injective.
@@
for any $|j|\leq i$, the operators $\{H_{i,l}\}_{i\geq 0}$ converges to an operator $H_{l}$ on $W_{\infty}$. In the Section A.4, we will prove that
@@
It follows from Propositions 2.15, 2.16, 2.17 and 2.19. ∎
@@ 2.7. The Boson-Fermion correspondence
$\rho(x_{i_{1}}\wedge x_{i_{2}}\cdots\wedge x_{i_{n}}):=y_{1}^{i_{n}-i_{n-1}-1}y_{2}^{i_{n-1}-i_{n-2}-1}\cdots y_{1}^{i_{1}-1}.$
@@
Now we consider the functor from the abelian category of vector spaces to the abelian category of admissible $\mathcal{H}$-representations
$\mathfrak{b}:W_{0}\to B\otimes_{\mathbb{Q}}W_{0}$
and the functor from the abelian category of vector spaces to the abelian category of graded admissible $\mathcal{E}$-representations
$\mathfrak{f}:W_{0}\to F\otimes_{\mathbb{Q}}W_{0}$
@@
The functors $\mathfrak{b}$ or $\mathfrak{f}$ is an equivalence between the category of vector spaces to the category of admissible $\mathcal{H}$-modules or the category of admissible $\mathcal{E}$ or $\mathcal{C}$-modules.
@@
For any admissible $\mathcal{H}$-module $W_{\infty}$, we have
$W_{\infty}\cong B\otimes_{\mathbb{Q}}W_{0}.$
@@
The morphism from $B\otimes_{\mathbb{Q}}W_{0}$ to $W_{\infty}$ by
$y_{1}^{l_{1}}\cdots y_{m}^{l_{m}}\otimes w_{0}\to H_{1}^{l_{1}}\cdots H_{m}^{l_{m}}w_{0}$
is a morphism of $\mathcal{H}$-modules. Now we prove that it is an isomorphism as vector spaces. By induction, we only need to prove that the morphism
$W_{i}\otimes_{\mathbb{Q}}\mathbb{Q}[y_{i+1}]$ $\to W_{i+1},$
$w_{0}+w_{1}y_{i+1}+\cdots+w_{n}y_{i+1}^{n}$ $\to w_{0}+H_{i+1}w_{1}+\cdots+H_{i+1}^{n}w_{1}.$
is an isomorphism. It is injective as given a monomial
$w=w_{0}+w_{1}y_{i+1}+\cdots+w_{n}y_{i+1}^{n}$
such that $w_{i}\in W_{i}$ and $w_{n}\neq 0$, we have $H_{-i-1}^{n}w=n!w_{n}$ and its image is non-zero.
To prove that it is surjective, let $w\in W_{i+1}$. We have
$w=\sum_{k=0}^{N}H_{i+1}^{k}((1-H_{i+1}\mathcal{L}_{i+1})\mathcal{L}_{i+1}^{k}w).$
for a sufficient large $N$ and for any $k\geq 0$, by 2.5
$(1-H_{i+1}\mathcal{L}_{i+1})\mathcal{L}_{i+1}^{k}w\in W_{i}.$
Hence the morphism is surjective. ∎
@@ 3. Derived categories of derived Grassmannians
Jiang [12, 13, 11] extended the definition of Grassmannians and flag varieties such that $\mathcal{V}$ can be any perfect complex. Particularly, when $\mathcal{V}$ is a tor-amplitude $[0,1]$ perfect complex, i.e. locally a two term complexes of locally free sheaves
@@
the derived Grassmannian is quasi-smooth (i.e. the locally complete intersection scheme in the derived algebraic geometry setting). Moreover, Jiang [13] strengthen the result of Toda [28] for the semi-orthogonal decomposition of derived Grassmannians by given the Fourier-Mukai kernels.
The concept of derived Grassmannians and flag varieties included many important incidence varieties in geometric representation theory. We will refer to [25] and [20] for several examples. In this paper, we will focus on the example in Section 5, i.e. the moduli space of perverse coherent sheaves on the blow-ups introduced by Nakajima-Yoshioka [22].
Now we review several main results of Jiang [12, 13, 11]. We will adapt the language of derived algebraic geometry and $\infty$-categories in this section. The relation between derived geometry and (classical) birational geometry can be described through the degeneracy theory, which we explained in a companion paper [29].
In this section, we will always assume that $X$ is a derived scheme and $\mathcal{E}$ is a Tor-amplitude $[0,1]$ perfect complex on $X$.
@@ 3.1. Derived Grassmannians of a Tor-amplitude $[0,1]$ complex
Remark 3.2.
Jiang’s definition requires that $d\geq 1$. Here we allow $d=0$ be defining
$Gr_{X}(\mathcal{E},0)\cong X.$
@@
Given a quasi-smooth derived scheme $X$ over $\mathbb{C}$, we say that $X$ is classical if it has trivial derived structure, i.e. $\pi_{0}(X)\cong X$ where $\pi_{0}(X)$ is the underlying scheme. By Corollary 2.11 of [1] (which we also discussed in the Appendix A of [30]), $X$ is classical if and only if
@@
Let $v_{d}$ be the universal quotient map over $Gr_{X}(\mathcal{E},d)$
$pr_{\mathcal{E},d}^{*}\mathcal{E}\xrightarrow{v_{d}}\mathcal{V}_{d}$
Proposition 3.6 (Theorem 4.18 of [11]).
The cotangent complex
$L_{Gr_{X}(\mathcal{E},d)/X}\cong\mathcal{V}_{d}^{\vee}\otimes fib(v_{d})$
where $fib$ denotes the fiber (i.e. cylinder) of a map.
Corollary 3.7.
The relative canonical bundle of $pr_{\mathcal{E},d}$ is
$K_{Gr_{X}(\mathcal{E},d)}:=det(L_{Gr_{X}(\mathcal{E},d)/X})\cong pr_{\mathcal{E},d}^{*}det(\mathcal{E})\otimes det(\mathcal{V}_{d})^{-1}.$
@@
(3.1) $f_{!}\cong f_{*}(-\otimes K_{Y/X})[vdim(f)],\quad f^{!}\cong(-\otimes K_{Y/X})\circ f^{*}[vdim(f)]$
where $vdim(f)$ is the rank of the cotangent complex $L_{f}$, which is a (locally) constant for a quasi-smooth morphism.
@@ 3.2. The incidence varieties
Given $(d_{+},d_{-})\in\mathbb{Z}_{\geq 0}^{2}$, the incidence scheme $Inc_{X}(\mathcal{E},d_{+},d_{-})$ can be defined in two ways, which were proven to be equivalent in [11]:
@@
    (3.2) $nc_{X}(\mathcal{E},d_{+},d_{-}):=Gr_{Gr_{X}(\mathcal{E}^{\vee}[1],d_{-})}(\mathcal{E}_{-}^{\vee}[1],d^{+}).$
@@
and we will also make it short as $r^{+},r^{-}$ respectively if it does not make any confusion.
As $r^{+}$ and $r^{-}$ are both quasi-smooth and proper, the functor
@@
Here $QCoh$ denotes the $\infty$-category of quasi-coherent sheaves and $Perf$ denotes the $\infty$-category of perfect complexes.
Now we compute the left and right adjoint functor of $\Phi_{d_{+},d_{-}}$. Let $e$ the rank of $\mathcal{E}$. First we notice that the relative virtual dimension of $r^{+}$ and $r^{-}$ is
$vdim(r^{+})=d_{-}(-e+d_{+}-d_{-}),\quad vdim(r^{-})=d_{+}(e+d_{-}-d_{+}).$
and the virtual dimension of the incidence variety $Inc_{X}(\mathcal{E},d_{+},d_{-})$ over $X$ is
$d_{-}(-e-d_{-})+d_{+}(e-d_{+})+d_{-}d_{+}.$
Let $\Psi_{d_{+},d_{-}}^{L}$ and $\Psi_{d_{+},d_{-}}^{R}$ be the left and adjoint functor of $\Phi_{d_{+},d_{-}}$. By the Grothendieck duality 3.1, we have
$\Psi_{d_{+},d_{-}}^{L}\cong(r^{-}\times r^{+})_{*}(K_{r^{-}}[vdim(r^{-})]),\quad\Psi_{d_{+},d_{-}}^{R}\cong(r^{-}\times r^{+})_{*}(K_{r^{+}}[vdim(r^{+})])$

### v3

@@ 2.2. Highest representation theory of $\mathcal{E}$ and Fermionic Fock space
The Verma module of the highest weight representation of weight $\lambda$ is denoted as $F_{\lambda,\mathcal{E}}$, and the highest weight vector is denoted as 00$\rangle$0 ⟩. We denote $F_{\mathcal{E}}:=F_{1,\mathcal{E}}$.
@@
$E\to\lambda E,\quad K_{1}\to\lambda K_{1},\quad F\to\lambda^{-1}F,\quad K_{-1}\to\lambda^{-1}K_{-1}.$
Under this auto equivalence, the representation $F_{\lambda,\mathcal{E}}$ is mapping to $F_{\mathcal{E}}$.
The Fermionic Fock space $F$ has the basis
@@
for any sequence $0<i_{1}<i_{2}<\cdots<i_{k},k\geq 0$. Through straightforward calculations, we find that
@@
There is a $\mathcal{E}$-representation on the Fermionic Fock space $F$ such that
@@
where we set $x_{0}:=0$. For the induced $\mathcal{C}$-representation on $F$, the action of $P_{i}$ is wedging $x_{i}$ and the action of $Q_{i}$ is $\frac{\partial}{\partial x_{i}}$.
@@
Let $|1\rangle$ be the highest weight vector of $F_{\mathcal{E}}$. Then the map
$|1\rangle\to 1$
@@
Now we prove that the Fermionic Fock space $F$ is irreducible as a $\mathcal{E}$-module. By associating all $x_{i}$ with degree $1$, we give a $\mathbb{Z}_{\geq 0}$-grading on $F$. Let $a=\oplus_{i=1}^{N}a_{i}\in F$ such that $a_{i}$ has degree $i$, we take a monomial
@@ 2.3. Admissible representations and Boson-Fermion correspondence
We consider the category of graded $\mathcal{E}$-modules (resp. $\mathcal{C}$-modules),
@@
A $\mathbb{Z}_{\geq 0}$-graded $\mathcal{E}$-module
@@
is called admissible if the action of $K_{-1}$ on $W_{i}$ is identity when $i=0$ and locally nilpotent when $i>0$.
Remark 2.9.
By definition we also have $K_{1}|_{W_{0}}=id$.
@@
    for any $x\in W$, $Q_{m}x=0$ when $m\gg 0$.
@@
On the Fermionic Fock space, we associate $x_{i}$ with degree $1$. Then $F$ is both a graded admissible $\mathcal{E}$-module and a graded admissible $\mathcal{C}$-module.
@@
$\begin{cases}H_{k}\to y_{k},\quad k>0,\\ H_{k}\to\frac{\partial}{\partial y_{-k}},\quad k<0.\end{cases}$
For a sequence of non-negative integers $\lambda=(n_{1},n_{2},\cdots)$ such that $n_{m}=0$ when $m\gg 0$, we denote $y_{\lambda}:=\prod_{i=1}^{\infty}y_{i}^{n_{i}}$, which is a monomial of finite variables.
@@
An $\mathcal{H}$-representation $W$ is called admissible, if for any $x\in W$ there exists an integer $N>0$ such that $H_{-\lambda}x=0$ for any $|\lambda|>N$.
Remark 2.14.
The $\mathcal{H}$-representation $B$ is admissible.
@@
There is an equivalence among the abelian category of admissible $\mathcal{H}$-modules, graded admissible $\mathcal{C}$-modules, and graded admissible $\mathcal{E}$-modules.
@@ 2.4. Graded $\mathcal{E}$-modules
First we prove that for any $x$ in $W$, $Q_{m}x=0$ when $m\gg 0$ by induction. If $x\in W_{0}$, it follows from the fact that $Q_{1}x=0$ and $Fx=0$. Assuming it is true for $x\in W_{i}$, then for $x\in W_{i+1}$, we have that for some $N>0$
$x=\sum_{i=1}^{N}E_{i}F_{i}x$
By A.5 and induction, we have
$Q_{m}E_{i}F_{i}x=E_{i}Q_{m-i}F_{i}x=0$
when $m\gg 0$. Hence $Q_{m}x=0$ for $m\gg 0$.
Now we prove that
@@
First we notice that for any non-zero $x\in\oplus_{i>0}W_{i}$, there exists $M\geq 0$ such $F_{M}x\neq 0$. Otherwise, by 2.2, for any $M>0$
@@
which contradicts the assumption of admissibility. Let $M_{0}$ be the smallest integer such that $F_{M_{0}}x\neq 0$. Then by induction, we have
$Q_{M_{0}+1}x=K_{1}Q_{M_{0}}K_{-1}x=\cdots=K_{1}^{M_{0}}Q_{1}K_{-1}^{M_{0}}x=K_{1}^{M_{0}+1}F_{M_{0}}x.$
@@
On the other hand, it is more complicated to construct a $\mathcal{E}$-representation from a graded $\mathcal{C}$-module, which we will leave to Proposition A.4.
@@ 2.5. Filtration and Fermionization
It satisfies that $W_{i}\subset W_{j}$ for $i\leq j$ and
@@
(2.5) $(Id-H_{k}\mathcal{L}_{k})W_{k}\subset W_{k-1}.$
Let $E:W_{k}\to W_{k+1}$ be the embedding operator, and $F:W_{k+1}\to W_{k}$ be the operator induced by $Id-H_{k}\mathcal{L}_{k}$. Moreover, we consider the operators
@@
where $H_{0}=\mathcal{L}_{0}:=id$. Then $W$ is an graded admissible $\mathcal{E}$-module under the action of $E,F,K_{1},K_{-1}$.
@@ 2.6. Stable limit and Bosonization
we consider the stable limit,
$W_{\infty}:=\lim_{n\to\infty}W_{i}$
where the morphism form $W_{i}$ to $W_{i+1}$ is $K_{1}$.
@@
for any $l\neq 0$, the operators $\{H_{i,l}\}_{i\geq 0}$ converges to an operator $H_{l}$ on $W_{\infty}$. In the Section A.4, we will prove that
@@
It follows from Propositions 2.16, A.4, 2.17 and 2.19. ∎
@@ 2.7. The Boson-Fermion correspondence
$\rho(x_{i_{1}}\wedge x_{i_{2}}\cdots\wedge x_{i_{n}}):=y_{1}^{i_{n}-i_{n-1}-1}y_{2}^{i_{n-1}-i_{n-2}-1}\cdots y_{n}^{i_{1}-1}.$
@@
By tensoring the Fermionic or Bosonic Fock space, it induces an equivalence from the abelian category of vector spaces to the abelian category of graded admissible $\mathcal{C}$-modules or admissible $\mathcal{H}$-modules.
@@
For any graded admissible $\mathcal{C}$-module $W$,
$W\cong F\otimes_{\mathbb{Q}}W_{0}.$
@@
For any $k>0$, we consider
$\bar{W}^{k}:=\bigcap_{l>k}ker(Q_{l})$
which is a $\mathbb{Z}_{\geq 0}$-grade vector space and
$W=\lim_{k\to\infty}\bar{W}^{k}.$
We notice that
$F\cong\lim_{k\to\infty}\wedge_{i=1}^{k}(\mathbb{Q}\oplus\mathbb{Q}x_{i}).$
Hence to prove Proposition 2.22, we only need to prove that for any $k\geq 0$,
$\bar{W}^{k}\cong W_{0}\otimes_{\mathbb{Q}}\bigotimes_{i=1}^{k}\wedge^{i}(\mathbb{Q}\oplus\mathbb{Q}x_{i}).$
Moreover, by induction, we only need to prove that
$\bar{W}^{k+1}\cong\bar{W}^{k}\otimes_{\mathbb{Q}}(\mathbb{Q}\oplus\mathbb{Q}x_{k+1}).$
as $\bar{W}^{0}\cong W_{0}$. It follows from the action of the Clifford algebra generated by $P_{k+1}$ and $Q_{k+1}$. ∎
@@ 3. Derived categories of derived Grassmannians
Jiang [13, 14, 12] extended the definition of Grassmannians and flag varieties such that $\mathcal{V}$ can be any perfect complex. Particularly, when $\mathcal{V}$ is a tor-amplitude $[0,1]$ perfect complex, i.e. locally a two-term complexes of locally free sheaves
@@
the derived Grassmannian is quasi-smooth (i.e. the locally complete intersection scheme in the derived algebraic geometry setting). Moreover, Jiang [14] strengthens the result of Toda [28] for the semi-orthogonal decomposition of derived Grassmannians by given Fourier-Mukai kernels. In this section, we review several main results of Jiang [13, 14, 12].
In this section, we will always assume that $X$ is a derived scheme and $\mathcal{E}$ is a Tor-amplitude $[0,1]$ perfect complex on $X$ of rank $e$.
@@ 3.1. Derived Grassmannians of a Tor-amplitude $[0,1]$ complex
Given a quasi-smooth derived scheme $X$ over $\mathbb{C}$, we say that $X$ is classical if it has a trivial derived structure, i.e. $\pi_{0}(X)\cong X$ where $\pi_{0}(X)$ is the underlying scheme. By Corollary 2.11 of [1] (which we also discussed in the Appendix A of [30]), $X$ is classical if and only if
@@
(3.1) $f_{!}\cong f_{*}(-\otimes K_{Y/X})[vdim(f)],\quad f^{!}\cong(-\otimes K_{Y/X})\circ f^{*}[vdim(f)].$
@@ 3.2. The incidence varieties
Given $(d_{+},d_{-})\in\mathbb{Z}_{\geq 0}^{2}$, the incidence scheme $Inc_{X}(\mathcal{E},d_{+},d_{-})$ can be defined in two ways, which are equivalent due to [12]:
@@
    $Inc_{X}(\mathcal{E},d_{+},d_{-}):=Gr_{Gr_{X}(\mathcal{E}^{\vee}[1],d_{-})}(\mathcal{E}_{-}^{\vee}[1],d_{+}).$
@@
which we will also abberivate as $r^{+},r^{-}$ respectively.
By Theorem 3.4, $r^{+}$ and $r^{-}$ are both quasi-smooth and proper, and the relative virtual dimensions are
$vdim(r^{+})=d_{-}(-e+d_{+}-d_{-}),\quad vdim(r^{-})=d_{+}(e+d_{-}-d_{+}).$
The virtual dimension of the incidence variety $Inc_{X}(\mathcal{E},d_{+},d_{-})$ over $X$ is
$d_{-}(-e-d_{-})+d_{+}(e-d_{+})+d_{-}d_{+}.$
Hence the functor
@@
Moreover, let $\Psi_{d_{+},d_{-}}^{L}$ and $\Psi_{d_{+},d_{-}}$ be the left and adjoint functor of $\Phi_{d_{+},d_{-}}$. By the Grothendieck duality 3.1, we have
$\Psi_{d_{+},d_{-}}^{L}\cong(r^{-}\times r^{+})_{*}(K_{r^{-}}[vdim(r^{-})]),\quad\Psi_{d_{+},d_{-}}\cong(r^{-}\times r^{+})_{*}(K_{r^{+}}[vdim(r^{+})])$
@@
Remark 3.6.
We will not calculate $K_{r^{-}}$ and $K_{r^{+}}$ explicitly in this paper as we do not need it. The reader can refer to [13] for the explicit formula. However, we should point out that they are generated by tautological complexes over the incidence variety.